11 11 TEX L^{A}TEX

PLIP Install HOWTO

Gilles Lamiral, lamiral@mail.dotcom.fr v1.15, 26 November 1999 This paper describes how to install a Gnu|Linux distribution on a computer without Ethernet card, nor cdrom, but just a local floppy drive and a remote nfs server attached by a Null-Modem parallel cable. The nfs server has a cdrom drive mounted and exported.

Contents

1 Introduction 1 2 License 2 3 How did I wrote this howto ? 4 4 What do you need ? 4 5 The documentation 5 6 Conventions 5 7 Network parameters 6 8 PLIP on the Source side 8 9 The cdrom nfs server side 11 10 FIPS, splitting of harddisk partitions 12 11 Debian installation 14

11.1 Preparing the two floppies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 11.2 The real installation process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 11.3 A install break: PLIP on the target side . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 11.4 Return to the normal install process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 11.5 Installing from a DOS partition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

12 Install the plip interface permanently 26 12.1 On the source side . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 12.2 On the target side . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

13 Need to compile a new kernel ? 29 14 What’s new 30 15 Todo 30 16 List of contributors and acknowledgements 30 You can find the last release of the PLIP-Install-HOWTO, by Gilles Lamiral, located at:

http://www.linuxdoc.org/HOWTO/PLIP-Install-HOWTO.html

1 Introduction

Writing a PLIP install HOWTO seems useless nowadays since Ethernet cards are cheap: NE2000 cards cost near the same as a Null-Modem cable. This is true for desktop computers, but not for laptop computers, which the PCMCIA card costs more than 10 times as the Null-Modem cable. Moreover, there is a parallel port on every computer, not a network card. Of course, this howto can be used to install Linux on every personal computer without loss of generality. This paper is just what I’ve done to install a Debian GNU/Linux distribution on a Toshiba Portégé 620CT laptop, from a NFS exported cdrom drive, via a Null-Modem cable. A Null-Modem cable is also called a LapLink cable but this word is trademarked by Traveling Software under the number 75466713 since 1986, so I won’t use it anymore.

2. License 2

This HOWTO will be obsolete when every Linux distribution include a PLIP install option. For example, the Debian installation only needs to add two commands to make this HOWTO obsolete (ifconfig + route). I hope one Debian maintainer will consider this point.

I would be happy to know if someone used the PLIP-Install-HOWTO to install other Linux distributions from other network protocols (ftp, http, nfs, samba, or even NT/Novell servers)

Feedback for typos, bad English, comments, money, job writings, joy, fears, cries are welcome and recommended (not with the same eagerness, choose yours).

2 License

I put this documentation under the opencontent <http://www.opencontent.org/opl.shtml> license. This means this document is given without any warranty. You can use it, redistribute it, modify it, by respecting the following conditions: You have to clearly specify the nature and the content of the modifications, their date and keep the opencontent license in case of redistribution. If you sell it, you sell the medium, not the content. This is a summary, read the license to get more details.

For those who can’t read the license online, I give a textual copy:

OpenContent License (OPL)
Version 1.0, July 14, 1998.
This document outlines  the  principles underlying the  OpenContent
(OC)  movement and may    be   redistributed provided it    remains
unaltered. For legal purposes,  this document is the license  under
which OpenContent is made available for use.
The  original   version  of this      document may  be   found   at
http://www.opencontent.org/opl.shtml
LICENSE
Terms and Conditions for Copying, Distributing, and Modifying
Items other than  copying, distributing, and modifying  the Content
with  which this license was distributed  (such as using, etc.) are
outside the scope of this license.
1.  You  may copy and distribute exact  replicas of the OpenContent
(OC)  as   you receive it,    in any  medium,   provided  that  you
conspicuously and appropriately publish on each copy an appropriate
copyright notice  and disclaimer of  warranty; keep  intact all the
notices  that refer  to this   License and to   the absence  of any
warranty; and give  any other recipients of the  OC a copy of  this
License along with the OC.  You may at your option charge a fee for
the media and/or handling involved in creating a unique copy of the
OC for use   off-line, you may at  your  option  offer instructional
support for the OC in exchange for a fee, or you may at your option
offer warranty in exchange for a fee. You  may not charge a fee for

2. License 3

the OC itself.  You may  not charge a fee  for the sole service  of
providing  access to and/or  use of the  OC via a network (e.g. the
Internet), whether it be via the world wide web,  FTP, or any other
method.
2.  You may modify your  copy or copies of  the  OpenContent or any
portion   of it,  thus   forming works  based on  the  Content, and
distribute such modifications or work  under the terms of Section 1
above, provided that you also meet all of these conditions:
a) You  must cause the  modified content to carry prominent notices
stating that you  changed it, the exact nature  and content  of the
changes, and the date of any change.
b) You must cause any work that you  distribute or publish, that in
whole or in part  contains or is  derived from the  OC or  any part
thereof, to be  licensed  as a  whole  at no  charge to  all  third
parties under the terms of this License, unless otherwise permitted
under applicable Fair Use law.
These requirements  apply  to the   modified  work as  a whole.  If
identifiable sections of that work are not derived from the OC, and
can  be reasonably considered   independent  and separate works  in
themselves, then this License, and its terms, do not apply to those
sections when  you distribute them as  separate works. But when you
distribute the same  sections as part of  a  whole which  is a work
based on the OC, the distribution of the whole must be on the terms
of this License,  whose permissions for  other licensees  extend to
the entire whole, and thus to each and every part regardless of who
wrote it.   Exceptions  are  made to  this  requirement to  release
modified works free of charge under this license only in compliance
with Fair Use law where applicable.
3. You are not required to accept this License,  since you have not
signed  it. However,  nothing else  grants you  permission to copy,
distribute or modify the OC. These actions are prohibited by law if
you  do not accept this    License.  Therefore, by distributing  or
translating the OC,  or  by deriving works herefrom,  you  indicate
your acceptance of this  License to do so,   and all its  terms and
conditions for copying, distributing or translating the OC.
NO WARRANTY
4.  BECAUSE THE OPENCONTENT (OC)  IS LICENSED FREE OF CHARGE, THERE
IS  NO WARRANTY FOR THE OC,  TO THE  EXTENT PERMITTED BY APPLICABLE
LAW. EXCEPT WHEN OTHERWISE  STATED IN WRITING THE COPYRIGHT HOLDERS
AND/OR OTHER PARTIES PROVIDE THE OC "AS IS" WITHOUT WARRANTY OF ANY
KIND, EITHER EXPRESSED OR IMPLIED,  INCLUDING, BUT NOT LIMITED  TO,
THE IMPLIED  WARRANTIES OF    MERCHANTABILITY  AND FITNESS  FOR   A
PARTICULAR PURPOSE. THE ENTIRE  RISK OF USE OF  THE OC IS WITH YOU.
SHOULD THE OC PROVE  FAULTY, INACCURATE, OR OTHERWISE  UNACCEPTABLE

          YOU ASSUME THE COST OF ALL NECESSARY REPAIR OR CORRECTION.
          5.  IN NO EVENT  UNLESS REQUIRED BY APPLICABLE LAW  OR AGREED TO IN
          WRITING  WILL ANY COPYRIGHT  HOLDER,  OR ANY   OTHER PARTY WHO  MAY
          MIRROR AND/OR REDISTRIBUTE THE OC  AS PERMITTED ABOVE, BE LIABLE TO
          YOU  FOR DAMAGES,  INCLUDING ANY  GENERAL,  SPECIAL,  INCIDENTAL OR
          CONSEQUENTIAL DAMAGES ARISING OUT OF   THE USE OR INABILITY TO  USE
          THE OC, EVEN IF SUCH HOLDER OR OTHER PARTY  HAS BEEN ADVISED OF THE
          POSSIBILITY OF SUCH DAMAGES.

3 How did I wrote this howto ?

I wrote this paper during the installation process but since I encountered several problems in the same time, my notes were disparate and sometimes I forgot to write the good command lines because I was fed up to reboot (kernel compilations) and edit this file each time.

When the installation went wrong (It did, I’m a beginner on laptops, plip, linux nfs, portmaper, Debian), I tried several things to fix the problems. When they remained too long, my brain was focused on solving, not writing.

So, when the installation process finally came out, I decided to restart the process from scratch by noting everything in order to save your time. It was the first time I installed Linux twice on the same computer.

Then, in order to verify the document, I again restarted from scratch with a Slink Debian distribution (The next after the Hamm one). It was the first time I installed Linux 3 times on the same computer. I hope it was the last time I do that.

Conclusion: If you do what is in this paper, it should work.

4 What do you need ?

• A Personal Computer, laptop or desktop, called the target computer or simply target.

• A 3"1/4 inch floppy drive on the target.

• Two or three fresh floppies. They don’t need to be formatted. You also need one more DOS formated floppy if you have to play with FIPS.

• A other Computer with a cdrom drive and nfs services, called source computer or simply source. It doesn’t have to be a Linux system but just a system which can export a cdrom drive via nfs. In this HOWTO, I assume it is a Linux system.

• A cdrom distribution. I took a Debian Hamm (2.0r3) from an old Infomagic compilation. I think every distribution can be installed this way, but I’m not sure. I’ve read a french document, written by Chmouel Boudjnah, saying the RedHat distribution has a easy installation process. The Chmouel’s document also deals with the Debian distribution, so if you read French, you can use the Chmouel’s document instead of this one.

• A Null-Modem DB 25 cable.

• Some time: 3 hours.

• Coffee.

5 The documentation

I recommend those good readings when you can’t figure out a problem. I’ve read them. You should too if you’re curious or conscientious.

Please, do not forget to consider the LDP mirrors, listed at:

<http://www.linuxdoc.org/mirrors.html>

Mostly every HOWTO are translated in many languages. Just go to a mirror bottom page and follow the

translations/ link.

If you want to use Linux on a laptop, read the last Laptop-HOWTO, by Werner Heuser, located at:

<http://www.snafu.de/~wehe/index_li.html>.

If you haven’t install any distribution yet, read carefully the Installation-HOWTO, by Eric S. Raymond, located at:

<http://www.linuxdoc.org/HOWTO/Installation-HOWTO.html>. And make more coffee :-)

If you need information about your cdrom, drive read the CDROM-HOWTO, by Jeff Tranter, located at:

<http://www.linuxdoc.org/HOWTO/CDROM-HOWTO.html>.

If you haven’t install a nfs server yet, read the NFS-HOWTO, by Nicolai Langfeldt, located at:

<http://www.linuxdoc.org/HOWTO/NFS-HOWTO.html>.

If you need to compile a new kernel on the source box, read the Kernel-HOWTO, by Brian Ward, located at:

<http://www.linuxdoc.org/HOWTO/Kernel-HOWTO.html>.

If you are new with plip, read the PLIP MINI-HOWTO, by Andrea Controzzi, located at:

<http://www.linuxdoc.org/HOWTO/mini/PLIP.html>.

If you plan to build your Null-Modem cable on your own, a good reading is the file PLIP.txt, by Donald Becker, located at:

/usr/src/linux/Documentation/networking/PLIP.txt

on your source linux computer, Luke.

6 Conventions

The file contents and line commands and install screen-shots are always in typewriter font, like this:

     #!/bin/bash
     #############################################
     #### This is the great file /bin/Windows ####
     #############################################
     while [ "1" ]; do
             echo "I do my best because I'm the best"

             echo "Very soon, next Y2Kill (the 01/01/0000)"
             echo "A new marvelous 64 bit release !"
             echo "Please wait a little more"
             sleep 18446744073709551615 # 2^64-1
     done

or this:

     $ killall Windows
     Terminated

The file content lines never begin with white space. You’ll have to remove them, if any. Sorry, I’m fed up with C-a M-AltGr-\ (remember, I’m a French azerty writer). Tab-emacs reflex is untamable.

Command input lines begin with a dollar $ (the prompt), you don’t have to type the dollar, just type the rest of the line; other lines are the command output, you don’t have to type them neither.

Because all the configuration commands are important, you’ll need to use a system administrator shell, like root, on the source and the target computers.

     $ su
     Password: blabla
     #

The prompt will remain "$" in this documentation, even if it should be "#". This is because "#" often means comment, so it is ambiguous. I don’t like ambiguity in computer science.

7 Network parameters

We will make a plip pointopoint network with those IP configuration:

• target: 192.168.0.1 netmask 255.255.255.255

• source: 192.168.0.2 netmask 255.255.255.255

You should not already have those names in your name space:

     $ ping source
     ping: unknown host source
     $ ping target
     ping: unknown host target

You should not already have those IP addresses in your network space:

$ ping  192.168.0.1
PING 192.168.0.1 (192.168.0.1): 56 data bytes
ping: sendto: Network is unreachable
ping: wrote 192.168.0.1 64 chars, ret=-1
--- 192.168.0.1 ping statistics ---
1 packets transmitted, 0 packets received, 100% packet loss
$ ping 192.168.0.2
PING 192.168.0.2 (192.168.0.2): 56 data bytes
ping: sendto: Network is unreachable
ping: wrote 192.168.0.2 64 chars, ret=-1
--- 192.168.0.2 ping statistics ---
1 packets transmitted, 0 packets received, 100% packet loss

If those commands don’t give you a error, change the names or the addresses.

You can choose other addresses, names or netmask (netmask must be the same on both sides). In the rest of this document, I’ll suppose you choose those addresses and those names.

Add a line in /etc/exports
#### file /etc/exports ####
...
/cdrom          source(ro) target(ro)
#### EOF ####

Because the portmaper do try to resolve IP addresses, add the IP addresses and names in /etc/hosts

#### file /etc/hosts ####
...
192.168.0.1     target
192.168.0.2     source
#### EOF ####

Verify you have the item files for the hosts search list in the file /etc/nsswitch.conf

#### file /etc/nsswitch.conf ####
...
hosts:      files nis dns
...
#### EOF ####

8 PLIP on the Source side

This section describes how to set the plip interface in the source server. If you run into troubles, I suggest to read the PLIP MINI-HOWTO.

Check if your lp device is not set. You should not have this entry:

$ cat /proc/devices
Character devices:
...
6 lp
...

If you have it, kill the lpd daemon and remove the lp module:

$ /etc/rc.d/init.d/lpd.init stop
Shutting down lpd: lpd
$ rmmod lp

If you can’t remove the lp module then you have to recompile the kernel with lp service as a module.

Now, the "6 lp" line is disappeared from the /proc/devices file, which is a reflection of the kernel possibilities.

You are not obliged to eliminate the lp device because it can work with it but it isn’t sure (it works for me). Check it yourself.

Check if your parallel port is handle:

$ ls /proc/parport/
0/
$ cat /proc/parport/0/hardware
base:   0x378
irq:    7
dma:    none
modes:  SPP,ECP,ECPEPP,ECPPS2

If you don’t have any directory under /proc/parport/ then you have to load the parport and parport_pc

modules:
$ insmod parport
$ insmod parport_pc

You should see this new entry in /var/log/messages:

Oct  9 20:50:47 louloutte kernel:
parport0: PC-style at 0x378 [SPP,ECP,ECPEPP,ECPPS2]
Oct  9 20:50:47 louloutte kernel:
parport0: detected irq 7;
use procfs to enable interrupt-driven operation.

I repeat the message "detected irq 7, use procfs to enable interrupt-driven operation", so:

$ echo 7 >  /proc/parport/0/irq

Check if plip module is loaded:

$ lsmod |grep plip

If plip module is not loaded, then load it:

$ insmod plip

You should see something like this in /var/log/messages

==> /var/log/messages <==
Oct  8 16:34:12 louloutte kernel:
NET3 PLIP version 2.3-parport gniibe@mri.co.jp
Oct  8 16:34:12 louloutte kernel:
plip0: Parallel port at 0x378, using IRQ 7

If you can’t load the plip module then you have to recompile the kernel with plip service as a module.

The syslog message says the module is loaded on the plip0 interface. Configure the plip0 interface:

$ ifconfig plip0 source pointopoint target netmask 255.255.255.255 up

Check if everything is ok.

$ ifconfig plip0
plip0     Link encap:10Mbps Ethernet  HWaddr FC:FC:C0:A8:00:02
     inet addr:192.168.0.2  P-t-P:192.168.0.1  Mask:255.255.255.255

     UP POINTOPOINT RUNNING NOARP  MTU:1500  Metric:1
     RX packets:0 errors:0 dropped:0 overruns:0
     TX packets:0 errors:0 dropped:0 overruns:0
     Interrupt:7 Base address:0x378

Now you can ping locally the source server:

$ ping source
PING source (192.168.0.2): 56 data bytes
64 bytes from 192.168.0.2: icmp_seq=0 ttl=255 time=0.3 ms
64 bytes from 192.168.0.2: icmp_seq=1 ttl=255 time=0.2 ms
--- source ping statistics ---
2 packets transmitted, 2 packets received, 0% packet loss
round-trip min/avg/max = 0.2/0.2/0.3 ms

Verify that the route to target exists:

$ route
Kernel IP routing table
Destination     Gateway         Genmask         Flags Metric Ref    Use Iface
target          *               255.255.255.255 UH    0      0        0 plip0

If the route doesn’t exist, add it:

$ route add -host  192.168.0.1  dev plip0

When the target will be configurated you can do a ping test:

$ ping target
PING 192.168.0.1 (192.168.0.1): 56 data bytes
64 bytes from 192.168.0.1: icmp_seq=0 ttl=64 time=4.5 ms
64 bytes from 192.168.0.1: icmp_seq=1 ttl=64 time=4.3 ms
--- 192.168.0.1 ping statistics ---
2 packets transmitted, 2 packets received, 0% packet loss
round-trip min/avg/max = 4.3/4.4/4.5 ms

But if you try it now you should have:

$ ping target

PING target (192.168.0.1): 56 data bytes
--- target ping statistics ---
5 packets transmitted, 0 packets received, 100% packet loss

Now, the server network is ready to work. Congratulations.

9 The cdrom nfs server side

This section describes how to mount and export via NFS a cdrom drive.

Mount the cdrom. If you encounter a problem with your cdrom drive, read the CDROM-HOWTO. I suppose that the cdrom device is /dev/hdd but it could be /dev/sr0 or /dev/hdb etc. The mount point I choose is /cdrom but you can choose the one you want:

$ mount -t iso9660 /dev/hdd  /cdrom

You have to set up nfs services. If something goes wrong, read the NFS-HOWTO

Verify your kernel supports nfs:

$ cat /proc/filesystems
        ext2
nodev   proc
nodev   nfs
nodev   ncpfs
nodev   devpts
        iso9660

Verify your portmaper can handle mountd and nfs client requests:

$ rpcinfo -p
program vers proto   port
100000    2   tcp    111  portmapper
100000    2   udp    111  portmapper
100005    1   udp    655  mountd
100005    1   tcp    657  mountd
100003    2   udp   2049  nfs
100003    2   tcp   2049  nfs

Lines with nfs or portmapper have to be there. NFS on tcp ? progress have been made !

Then, rerun portmap, mountd, nfs:

$ /etc/rc.d/init.d/portmap.init stop
Stopping INET services: portmap
$ /etc/rc.d/init.d/portmap.init start
Starting portmapper: portmap
$ /etc/rc.d/init.d/nfs stop
Shutting down NFS services: rpc.mountd rpc.nfsd
$ /etc/rc.d/init.d/nfs start
Starting NFS services: rpc.mountd rpc.nfsd

Then, try to mount it "locally" (via nfs) on the source box and clean the test:

$ mkdir /tmp/nfstest
$ mount -t nfs source:/cdrom /tmp/nfstest
$ ls  /tmp/nfstest/
README             debian/            locatedb.3         tools/
TRANS.TBL          ftp.netscape.com/  ls_lR.3            upgrade/
boot/              install/           realaudio/
$ echo great stuff !
$ umount /tmp/nfstest
$ rmdir /tmp/nfstest

Bad luck ? Read the NFS-HOWTO and /var/log/messages.

Check the nfs server with those commands:

$ rpcinfo -p

Run portmap with the -v flag:

 $ portmap -v

kill portmap, mountd, nfsd and rerun them in this order: portmap, mountd, nfsd. Take a rest. Read again the NFS-HOWTO and restart.

Now, the exported cdrom drive nfs server is ready to work. Congratulations.

10 FIPS, splitting of harddisk partitions

This section is for people who has a one only primary FAT partition (DOS, Windows 3.xx, NT) and wants to keep it without loosing data. Make some backup because if you burn it, I won’t be responsible. You will.

Download FIPS <ftp://ftp.debian.org/pub/debian/tools/>. Don’t forget the mirrors ftp.xx.debian.org where xx is your country abbreviation (fr, fi, us, uk, etc.).

At the time of this writing the archive file is called fips20.zip.

The fips program is already in the debian cdrom distribution. I found it (yes, after download the fips20.zip file) in /cdrom/debian/tools/fips15.zip (It supposes your cdrom is mounted on /cdrom)

I guess you are on a Unix world but you’re not compelled to do so. Go in a good working place on the source computer:

     $ mkdir /tmp/fips-2.0/
     $ cd /tmp/fips-2.0/
     $ unzip -l /archive/fips/fips20.zip
     ...
     $ unzip  /archive/fips/fips20.zip
     ...
     $ ls
     $ dos2unix fips.doc fips.faq  readme.1st

1. Read the file readme.1st

2. Read the file fips.doc

3. Read the file fips.faq

Boot your target computer. Be in pure DOS (Quit Windows).

Read again the file fips.doc from the section "5. Before you start".

Be aware of hidden files.

Be aware to eliminate the "virtual memory" file (swap for Win*) during the fips process. In Windows 3.11 (Quite up to date, no?) this swap file is configurated from Program-Manager->Control-Panel->Extent (a i386 chip icon).

Click->Click->Click on the chip and re-click on a button called "Virtual Memory".

Adjust the size to none, Click<-Click<-Click<-Click back plus Alt-F4 to close all your windows. Don’t you think the gates are too closed too ?

Go to sleep.

Run SCANDISK:
     C:\> SCANDISK

Correct every cluster. You should have no dead cluster now.

Quit SCANDISK

Run DEFRAG on C:

     C:\> DEFRAG C:

All clusters are at the beginning of the harddisk.

Quit DEFRAG

Make a bootable floppy disk:

     C:\> FORMAT A:/S

Copy AUTOEXEC.BAT and CONFIG.SYS on the new floppy if you want to keep your local features (keyboard keys etc.). Remove the line running SMARTDRIVE.EXE in AUTOEXEC.BAT. Keep a minimal AUTOEXEC.BAT

On the source box: Copy the fips files restorrb.exe, fips.exe and errors.txt to this floppy disk.

    $ mcopy errors.txt fips.exe restorrb.exe a:

Boot the target with the floppy. Run fips while reading the section "Using FIPS" in the file fips.doc

Do not forget to answer yes when fips ask for a rootboot saving.

When everything is ok, save your hard disk space:

     $ cd
     $ rm -rf /tmp/fips-2.0/

Now you have a nice free space for a new operating system.

11 Debian installation

I’m not a Debian proselyte since it is my first Debian installation. Debian lovers are the same kind of people as Macintosh lovers or Linux lovers in the operating systems space. Nothing else counts. Since I’m already a Mac and Lin lover (and French too :-) it was time to fall in love.

Debian is well, clearly and internationally documented. Thanks to all those guys who bred this really open distribution.

I could leave you on your own during the installation process. But since we have to interact with a shell during it, the entire process will be described in details.

I know the description is Debian specific. I prefer to give you a complete example than nothing except a “run a shell at the right moment and type bla bla...”. I do like concrete examples.

Simon Forget <sforget@camelot.ca> told me he could not use this howto with his Toshiba Libretto 50CT because the kernel could not recognize the pcmcia floppy drive during the installation process. I don’t know why but there is a solution with plip, if you already have a dos partition on the target computer. This solution is simpler and faster because no floppy drive nor plip network are necessary during the installation process. If you are interested by this solution, go directly to the section 11.5 (Installing from a DOS partition). I keep the old one installation process because this new one needs a DOS partition that becomes unnecessary when you definitively want to leave the dark side.

11.1 Preparing the two floppies

On the source side, mount the cdrom and go in the install directory

      $ mount -t iso9660 /dev/hdd  /cdrom
      $ cd /cdrom/debian/dists/stable/main/disks-i386/current/

Read the file install.html with a browser or install.txt with a cat, less or more (a dog). Now, record the install (alias rescue) floppy. Write a "resc1440.bin" label on it:

      $ dd if=resc1440.bin of=/dev/fd0H1440

Record the drivers floppy. Write a "drv1440.bin" label on it:

      $ dd if=drv1440.bin of=/dev/fd0H1440

Lock the writing on those floppies. Now you’re ready to start the real installation process.

11.2 The real installation process

Insert the resc1440.bin floppy on target drive. Reboot your target box.

     Welcome to Debian GNU/Linux 2.x!
     ...

Read the text The prompt is :

     boot:

Press <ENTER>

     Loading root.bin...........
     loading linux...
     ....

A new screen:

     Next: Select Color or Monochrome display

Choose yours with the arrow up/down keys and press <ENTER>

     Next: Continue with the installation

press <ENTER>

     Software in the Public Interest
             presents
      *** Debian GNU/Linux 2.1 ***
     ...

Read the text Press <ENTER> (You see <Continue> on the screen)

     Next: Configure the Keyboard

press <ENTER> Select your country and press <ENTER>

     Next : Partition a Hard Disk

Press <ENTER>

     Select Disk Drive
     /dev/hda

If you have several disks choose the good one (where you want to install Linux) and Press <ENTER> You are now using the program cfdisk. Using cfdisk is safe until you decide to [Write] the partition on disk. Leaving cfdisk with [Quit] is safe. We are going to make just two new primary partitions, a Swap one and a big Linux one. If you want to do more partitions, leave some free space or extended partitions, you can. But read the documentation about Partitioning in the Installation-HOWTO. If you previously ran the FIPS program, you see the hda1 partition (DOSFAT16). Switch to the next free partition with the up/down arrow keys. Select [New] with the right/left arrow keys and press <ENTER> Select [Primary] and press <ENTER> Enter the size of your swap partition. Twice the ram is usual if you have less than 128 mega bytes of ram. If you have 2 giga bytes of ram, it’s because you don’t want to swap. In that case, no swap partition is needed.

Select [Beginning] and press <ENTER>

Select [Type] and press <ENTER> Type 82 (Linux Swap) and press <ENTER> Switch to the next free partition with the up/down arrow keys Select [New] and press <ENTER> Select [Primary] and press <ENTER> Enter the size in MB (You can leave the default) and press <ENTER> Its type is already Linux. If not, change to Linux (83) with [Type] The partition table is defined now. Verify everything looks good. If you’re not sure, read the documentation with [Help]. If doubts are still there, select [Quit] and leave the installation process. Go for a walk and restart from the beginning of this section. I assume your are confident now.

Select [Write]
     Are you sure you want to write the partition table to disk?

Type yes and press <ENTER> Select [Quit] and press <ENTER>

     Next: Initialize and Activate a Swap Partition

Press <ENTER>

     Please select the partition to initialize as a swap device

Select /dev/hda2 (normally already selected) Press <ENTER>

     Scan for Bad Blocks?

select <YES> and press <ENTER>

     Are you Sure?

select <YES> and press <ENTER>

     Initializing swap partition
     ...
     Next: Initialize a Linux Partition

Press <ENTER>

     Select Partition Please select the to initialize as a Linux "ext2"
     file-system.

Select /dev/hda3 (normals already selected) and press <ENTER>

     Scan for Bad Blocks?

Select <YES> and press <ENTER>

     Are you Sure?

Select <YES> and press <ENTER> A new page full of numbers. You can take a rest because it takes some time (especially with big hard disk).

     Next: Mount a Previously-Initialized Partition

Press <ENTER>

     Please select the partition to mount

Select /dev/hda3 and press <ENTER>

     Mount the /dev/hda3 device as the Root FileSystem?

Select <Yes< and press <ENTER>

     Next: Install Operating System Kernel and Modules

Press <ENTER>

     Please select the medium you will use to install the system

Select /dev/fd0 and press <ENTER>

     Please place the Rescue Floppy in the first floppy drive

The floppy is already there. Select <Continue> and press <ENTER>

     Installing the Rescue Floppy ...
     Please place the Drivers Floppy in the first floppy drive

Eject the Rescue Floppy and insert the Drivers Floppy, the one I you labelled drv1440.bin (you did it, didn’t you?). Select <Continue> and press <ENTER>

     Installing the Drivers  Floppy ...

A new screen:

     Next: Configure Device Driver Modules

Press <ENTER>

     Select Category

Read the text.

     Please select the category of modules

Select net and press <ENTER> Select plip.

     Module plip

Select "Install the module in the kernel" and press <ENTER> No parameters are needed. Select <Ok> and press <ENTER>

     Installation succeeded
     Please press ENTER when you are ready to continue.

Press <ENTER> Select Exit (Finished with these modules) and press <ENTER> Then you see the same kind of screen again. Select Exit (Finished with these modules) and press <ENTER>

     Next: Configure the Network

Choose a name, you can use a different name than debian or target. Press <ENTER>

     Is your system connect to a network?

Select <No> and press <ENTER> STOP STOP STOP. You see now:

     Next: Install the base system

Now we need a shell. Press Alt F2 and <ENTER> You are in a root shell.

11.3 A install break: PLIP on the target side

You are in a root shell. Verify the plip module is loaded:

$ lsmod
Module    Pages    Used by
plip          3          0

Find the exact name of the plip interface:

$ dmesg
...
NET3 PLIP version 2.2 gniibe@mri.co.jp
plip1: Parallel port at 0x378, using assigned IRQ 7

You can find the same information with

$ cat /proc/kmsg
...
<4>NET3 PLIP version 2.2 gniibe@mri.co.jp
<4>plip1: Parallel port at 0x378, using assigned IRQ 7
...
Cntrl-c

Configure the plip interface:

$ ifconfig plip1 192.168.0.1  pointopoint  192.168.0.2 netmask 255.255.255.255 up

Verify everything is ok:

$ ifconfig plip1
plip0     Link encap:Ethernet  HWaddr FC:FC:C0:A8:00:01
     inet addr:192.168.0.1  P-t-P:192.168.0.2  Mask:255.255.255.255
     UP POINTOPOINT RUNNING NOARP  MTU:1500  Metric:1
     RX packets:0 errors:0 dropped:0 overruns:0 frame:0
     TX packets:0 errors:0 dropped:0 overruns:0 carrier:0
     Collisions:0
     Interrupt:7 Base address:0x378

Verify the route to source exists:

$ route
Kernel IP routing table
Destination     Gateway         Genmask         Flags Metric Ref    Use Iface
127.0.0.0       0.0.0.0         255.0.0.0       U     0      0        0 lo

If not, like in the previous example, add the route to 192.168.0.2, the source:

$ route add -host 192.168.0.2 dev plip1

Now the route is installed:

$ route
Kernel IP routing table
Destination     Gateway         Genmask         Flags Metric Ref    Use Iface
192.168.0.2     0.0.0.0         255.255.255.255 UH    0      0        0 plip1
127.0.0.0       0.0.0.0         255.0.0.0       U     0      0        0 lo

Now you can perform a successful ping from the source server (ping is not available on the Debian install process):

$ ping target
PING target (192.168.0.1): 56 data bytes
64 bytes from 192.168.0.1: icmp_seq=0 ttl=64 time=14.0 ms
64 bytes from 192.168.0.1: icmp_seq=1 ttl=64 time=4.3 ms
--- target ping statistics ---
2 packets transmitted, 2 packets received, 0% packet loss
round-trip min/avg/max = 4.3/9.1/14.0 ms

If it doesn’t work, check the Null-Modem cable connection, rerun dmesg, ifconfig, route. Verify everything. Take a rest. Restart. When the ping is ok, return to the normal installation process: On the target: Press Alt F1

11.4 Return to the normal install process

You are back to the normal installation process. You should still see:

Next: Install the base system

Press <ENTER>

Please select the medium you will use to install the system

Select nfs and press <ENTER>

Please choose the NFS server and the mount path ...
Type: 192.168.0.2:/cdrom

Press <ENTER>

Please choose the path inside the mounted NFS filesystem

Type /debian (normally already there) and press <ENTER> If no error message complains about the nfs mount then bravo. You can verify this mount by entering the shell again:

Press Alt F2
$ mount
...
192.168.0.2:/cdrom on /instmnt type nfs (rw, addr=192.168.0.2)
$ ls /instmnt/debian
README                 README.non-US          doc/
README.CD-manufacture  README.pgp@            hamm/
README.mirrors.html    TRANS.TBL              tools/
README.mirrors.txt     dists/
Press Alt F1

You’re back again to the normal installation process. Now the plip nfs cdrom connection is done. Let’s go on and finish our job.

Please select the directory containing a file base2_0.tgz

Select list and press <ENTER>

Please Wait
The installation program is building a list of ...

A new screen:

Select Archive Path
Please select the directory that you will use to install the Base
System from.

Only one long item, already selected. Press <ENTER>

The Base System is being extracted from
/instmnt/debian/dists/.....

Take a second rest, you deserve it

Next: Configure the Base System

Press <ENTER>

Select Timezone

Select your timezone and your directory and press <ENTER>

     Timezone Configuration

Read and press <ENTER>

     Timezone Configuration.

An other page Read, select <YES> and press <ENTER>

     Next: Make Linux Bootable Directly From Hard Disk

Press <ENTER>

     Create Master Boot Record?

Read Select <Yes> and press <ENTER>

     Make Linux the Default Boot Partition?

Read. Select <No> and press <ENTER>

     Next: Make a boot Floppy

Press <ENTER>

     Change Disk
     Please place a blank floppy disk in the first floppy drive.

Do it and press <ENTER> The floppy is being formated

     Creating a filesystem on the floppy...
     Copying the operating system kernel...

A new screen:

     Next: Reboot the System

Press <ENTER>

     Reboot the system ?

Remove the floppy and press <ENTER> The system reboot. Are you still with the dark side? I guess yes, so insert the boot floppy you’ve just made and reboot again with Ctrl-Alt-Del See the boot messages. Read the text.

     New password:

Enter a root password.

     Re-enter new password

Do it.

     Shall I create a normal user account now? [Y/n]

Enter n and press <ENTER>

     Shall I install shadow passwords? [Y/n]

Enter y and press <ENTER>

     Do you want to use a PPP connection to install

Enter n and press <ENTER>

     Now you may choose one of several selections ...
     Do you want to perform this step?

Enter n and press <ENTER>

     I'm going to start the 'dselect' program...

Press <ENTER> Select [Q]uit and press <ENTER>

     You may now login as 'root' at the login: prompt...
     ...
     debian login:

Enter root.

     Password:

Enter the root password.

    ...
     debian:~#

The system is installed and working. CONGRATULATIONS!

My job stops here. Read the install documentation of your distribution and go on with the Unix system administration job.

When you reboot your system, the plip connection won’t be in good shape. But now you now what to do.

I suggest you some work:

• Configure the plip interface for the normal boot process (in the file /etc/init.d/network).

• Configure /etc/fstat on the target to mount simply the remote source cdrom via nfs.

           source:/cdrom              /cdrom   nfs     noauto,intr 1 2

• Learn Lilo. Configure it for your DOS and Linux systems and install it on a floppy.

• When you master Lilo on the floppy, install it on your hard drive.

• Install and configure the X Window System.

• Have fun.

• Mail me a feedback.

11.5 Installing from a DOS partition

You can use this section if you already have a DOS partition on your target computer. I think it is a faster installation method.

Read the section 5.3.1 "Installing from a DOS partition" from install.txt. This section is also available on the Debian CD in the html file ch-install-methods.html.

I just going to help you to make the first point : "1. Get the following files from your nearest Debian []" repository.

The transfer is easy with tomsrtbt.

Export the source cdrom drive via nfs.

Read the file tomsrtbt.FAQ

Create the floppy under DOS or Linux, you have the choice.

     Linux installation:
     a) extract the .tar.gz archive
     b) Be root
     c) Be in the tomsrtbt-<version> directory
     d) Have a blank floppy with no bad sectors
     e) Do './install.s'

Boot the target with the floppy. Log in root.

      $ insmod plip
      $ ifconfig plip1 192.168.0.1 pointopoint 192.168.0.2 \
      netmask 255.255.255.255 up
      $ route add -host 192.168.0.2 dev plip1
      $ mount -t nfs 192.168.0.2:/cdrom /cdrom
      $ mkdir /c
      $ mount -t msdos /dev/hda1 /c
      $ mkdir /c/debian
      $ cd /cdrom/dists/stable/main/disks-i386/current/
      $ cp resc1440.bin drv1440.bin base2_1.tgz root.bin linux \
      install.bat loadlin.exe /c/debian

The \ means the command line continues on the next line. The cp command takes some time. When finished, remove the floppy and then :

      $ reboot

Boot under DOS.

      C:\> cd debian
      C:\DEBIAN> install

Here you go to install Debian from a DOS partition. You do not need plip nor floppy during the installation process. Everything deals with the hard drive. Read the Debian install documentation, it is a very good and clear one.

12 Install the plip interface permanently

12.1 On the source side

I use an old Linux RedHat 4.1 distribution. The location of the files can be different on other GNU/Linux distributions but the philosophy is the same (The Unix System V convention).

Create the file /etc/rc.d/init.d/plip with this content:

#!/bin/sh
##############################
# file /etc/rc.d/init.d/plip #
##############################

# See how we were called.
case "$1" in
start)
# Start daemons.
/bin/echo "Starting plip interface: "
/bin/echo "Doing /sbin/ifconfig plip0 source pointopoint target netmask 255.255.255.255 up"
/sbin/ifconfig plip0 source pointopoint target netmask 255.255.255.255 up
/bin/echo  "Doing /bin/ping -q -c 4 target"
/bin/ping -q -c 4 target
/bin/echo "Starting plip interface: done"
;;
stop)
# Stop daemons.
/bin/echo  "Shutting down plip interface:"
/bin/echo  "Doing /sbin/ifconfig plip0 source pointopoint target netmask 255.255.255.255 down"
/sbin/ifconfig plip0 source pointopoint target netmask 255.255.255.255 down
/bin/echo  "Doing /sbin/modprobe  -r plip "
/sbin/modprobe  -r plip
/bin/echo "Shutting down plip interface: done"
;;
*)
echo "Usage: $0 {start|stop}"
exit 1
esac
exit 0
# === End of File ===

Only the ifconfig lines are strictly necessary. Perhaps you will need to add some modprobe commands if you don’t use kerneld nor the kmod feature of new kernels 2.2.x

Create the symbolic links in the rc*.d directories:

$ cd /etc/rc.d/rc0.d/
$ ln -s ../init.d/plip K97plip
$ cd /etc/rc.d/rc1.d/
$ ln -s ../init.d/plip K92plip
$ cd /etc/rc.d/rc3.d/
$ ln -s ../init.d/plip S11plip

$ cd /etc/rc.d/rc5.d/

$ ln -s ../init.d/plip S11plip

You can choose other numbers. Be aware to place the K??plip files after the ones that shutdown services depending on plip. Be aware to place the S??plip files before the ones that start services depending on plip, nfs, nis, ftp, http etc.

Update the /etc/conf.modules file:
# /etc/conf.modules
...
alias parport_lowlevel parport_pc
post-install parport_pc echo 7 >  /proc/parport/0/irq
...

Choose the good irq number (7 is mine, not yours).

Test the plip shell:

$ /etc/rc.d/init.d/plip
Usage: /etc/rc.d/init.d/plip {start|stop}
$ /etc/rc.d/init.d/plip stop
Shutting down plip interface:
Doing /sbin/ifconfig plip0 source pointopoint target netmask 255.255.255.255 down
Doing /sbin/modprobe  -r plip
Shutting down plip interface: done
$ /etc/rc.d/init.d/plip start
Starting plip interface:
Doing /sbin/ifconfig plip0 source pointopoint target netmask 255.255.255.255 up
Doing /bin/ping -q -c 4 target
PING target (192.168.0.1): 56 data bytes
--- target ping statistics ---
4 packets transmitted, 4 packets received, 0% packet loss
round-trip min/avg/max = 4.4/8.3/14.0 ms
Starting plip interface: done

Updating the start scripts is a good accasion to reboot a Unix system, to check the modifications. Do it:

$ init 6

12.2 On the target side

Update the file /etc/init.d/network:

#! /bin/sh
#######################
# /etc/init.d/network #
#######################
ifconfig lo 127.0.0.1
route add -net 127.0.0.0
ifconfig plip1 192.168.0.1 pointopoint 192.168.0.2 netmask 255.255.255.255 up
route add -host 192.168.0.2 dev plip1

That’s all because the parport features are directly in the kernel.

Updating the start scripts is a good occasion to reboot a Unix system, to check the modifications. Do it:

$ init 6

13 Need to compile a new kernel ?

Several times during the source server configuration we mentioned that a kernel compilation on the source computer could be needed.

I list the things you need to change or set. If you have never compile a kernel, read the Kernel-HOWTO.

Loadable module support  --->
[*] Enable loadable module support
General setup  --->
[*] Networking support
<M> Parallel port support
<M>    PC-style hardware
Networking options  --->
[*] TCP/IP networking
Network device support  --->
[*] Network device support
<*> Dummy net driver support
<M> PLIP (parallel port) support
< > or <M>  PPP (point-to-point) support
< > or <M> SLIP (serial line) support

Character devices  --->
<M> Parallel printer support

Filesystems  --->
   [*] /proc filesystem support
   <*> or <M> ISO 9660 CDROM filesystem support
   Network File Systems  --->
       <*> or <M> NFS filesystem support

14 What’s new

v1.15, 26 November 1999

• The section "What’s new". It is this section.

• The section "Install the plip interface permanently".

• The section "List of contributors and acknowledgements".

• Made the different chapters dealing with the detailed Debian installation be a single chapter with the actual chapters as just subsections.

• Made a new section "Installing from a DOS partition". A simpler and faster method.

15 Todo

• Check a kernel compilation from scratch (no last .config file) with only the options I gave. Add the missing ones if any.

• Make the Red Hat, Mandrake, Caldera OpenLinux, SuSE, Turbo Linux, Slackware, Stampede, detailed installation chapters. Contributions are welcome. Gilles Lamiral won’t do that job unless given new computers and CDs.

• Make "The cdrom nfs server side" chapter become simply "the server side" chapter with subsections like "cdrom medium", "zip medium", "hard disk medium", and "ftp server" "nfs server", "samba server" "http server". Again, contributions are welcome.

16 List of contributors and acknowledgements

I thank all the people in this list of contributors or projects. If you think I forgot someone, do not hesitate to complain.

Werner Heuser <wehe@snafu.de> He said: "yesterday I had time enough to read your document. It’s great and in the next issue of my HOWTO (probably 2. December) I will make a link to it." LDP <linux-howto@metalab.unc.edu>

The Linux Documentation Project accepted the PLIP-Install-HOWTO without any hesitation.

Simon Forget <sforget@camelot.ca> Simon had a problem with his pcmcia floppy drive. So he made me search a solution that is, in fact, simpler and faster.